OpenStudy (anonymous):

plisss plissss help me "n" is the smallest positive integer with the condition "n" is divisible by 20, n ^ 2 is a cubic number, n ^ 3 is a square number. n + 1234567 =

OpenStudy (turingtest):

n=1,000,000=10^6 (10^6)/20=5(10^4) (10^6)^2=10^12=(10^4)^3 (10^6)^3=10^18=(10^9)^2 n+1234567=2234567

OpenStudy (anonymous):

owh thanks ^_^

OpenStudy (anonymous):

excuse me, I want to ask how to get (10^6)/20=5(10^4)???????? pliss explain to me :(

OpenStudy (turingtest):

ok, I used scientific notation there so I'm sorry if it threw you off. 10^6=1,000,000 1,000,000/20=50,000=5x10^4 in scientific notation we can make a number very large or small by multiplying it by ten to some power, since the exponent of ten tells you how many zeros you have after the number 10^3=1,000 so if I want to write the number 5,300 in scientific notation I write 5.3x10^3 because 5.3(1000)=5300

OpenStudy (anonymous):

owh excellent thanks ^_^

OpenStudy (turingtest):

no problem